17 11/12 diveded by 6 1/10 =
HELP PLEASE

Found 2 solutions by checkley71, brandombe:
Answer by checkley71(8403)

FIRST WE NEED TO CHANGE THE MIXED FRACTION TO AN IMPROPER FRACTION.
(17 11/12=(12*17+11)/12=215/12
NOW CHANGE THE OTHER FRACTION ALSO
6 1/10=(10*6+1)/10=61/10
NOW WE CAN DIVIDE THESE 2 FRACTIONS THUS:
(215/12)/(61/10) NOW WE NEED TO INVERT THE DENOMINATOR & MULTIPLY
215/12*10/61
2150/732
2.937 ANSWER.
----------------------------------------------------------------
OR WE COULD CHANGE THESE FRACTIONS TO DECIMALS AND DIVIDE.
17 11/12=17.9167
6 1/10=6.1
NOW DIVIDE:
17.9167/6.1
2.937 ANSWER.

Answer by brandombe(4)

17 11/12 divied by 6 1/10= 2 343/366